# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v1

@@ III-C Almost Necessary and Sufficient Conditions
In this section, we propose approximate remedies for non-minimum phase systems. Our goal is to relax, at the cost of losing exactness, the and condition in Theorem 1, $\bigl{(}Y(z)=0,\ \forall z\in Z_{R}\bigl{)}$. This can be achieved using the $\mathcal{H}_{\infty}$-Synthesis framework.
We define a virtual loop in the sense of classical feedback structure shown by Fig.1.(a). Given the system $P(s)$ as in (11), suppose we have a virtual controller, say $K(s)\in\mathbb{R}(s)^{n_{u}\times n_{y}}$, be defined such that $u_{\upsilon}(t)=K(s)(r_{\upsilon}(t)-y_{\upsilon}(t))$. The key transfer functions within the virtual feedback loop are defined as follows: the input complementary sensitivity function, $T_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}K(s)P(s)$; the input sensitivity function, $S_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}$; and the loop transfer function, $L_{i}(s)\triangleq K(s)P(s)$. By breaking the loop at the output gives the output functions, $T_{o},S_{o},L_{o}$.
@@
Lemma 1 (Youla’s Statements[26]).
Assume (28) holds. Then the following statements are true:
  1. 1.
    If $P(s)$ has a RHP zero at $z$, then $PK$,$KP$, $PK(I+PK)^{-1}$, $(I+PK)^{-1}P$, and $KP(I+KP)^{-1}$ each have a RHP-zero at $z$.
  2. 2.
    If $P(s)$ has a RHP-pole at $p$, then $PK$ and $KP$ also have a RHP-pole at $p$, while $(I+PK)^{-1}$, $K(I+PK)^{-1}$,$(I+KP)^{-1}$ and $(I+KP)^{-1}K$ each have a RHP-zero at $p$.
@@
$\begin{bmatrix}z(t)\\ e_{\upsilon}(t)\end{bmatrix}={\begin{bmatrix}W_{P}&W_{P}P\\ -I&-P\end{bmatrix}}\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}{\color[rgb]{0,0,0}=G\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}}$ (29)
where each scalar in $W_{P}$ is as defined in [33, Equation 22], $e_{\upsilon}(t)\triangleq r_{\upsilon}(t)-y_{\upsilon}(t)$, and $z(t)\triangleq W_{P}e(t)$. The LFT is then given by $\mathcal{F}_{l}(G,K)\triangleq{G}_{11}+{G}_{12}{K}({I}-{G}_{22}{K})^{-1}{G}_{21}$ where the $\mathcal{H}_{\infty}$ control problem involves finding $K$ such that
@@
By letting $\epsilon_{i}\to 0$, we replace all approximate integrators in $K$ with pure integrators [36] that imposes: If $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))<\mathbb{E}[N^{2}]$, then $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))=0$ where $\mathbb{E}[N^{2}]$ denotes the expected value mean square of random noise in dB .
@@
Consider $P(s)$ as defined in (11)-(12) with $\text{rank}_{\mathbb{R}(s)}(P)=r=n_{u}=n_{y}$, and $Z_{R}\neq\emptyset$. Define the approximate inverse $\Xi_{a}^{\ell}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$ as
$\Xi_{a}^{\ell}(s)=S_{i}(s)K(s),$ (31)
@@
$\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\quad\forall t>0,$
for some $\alpha,\beta>0$.
@@
As in Corollary 1, since $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, we have a unique solution $P^{-1}$ that yields $e(t)=0,\forall t>0$ but $\text{lim}_{t\to\infty}u_{inv}(t)=\infty$. Therefore to obtain a stable control input, we begin by using (10) and taking the Laplace transform:
@@
Next, substitute $U_{inv}(s)=S_{i}(s)K(s)Y(s)$ in (33), noting that $S_{i}(s)=(I+K(s)P(s))^{-1}$:
$E(s)=Y(s)-P(s)(I+K(s)P(s))^{-1}K(s)Y(s).$
Rewriting,
$E(s)$ $=Y(s)[I-\underbrace{(I+P(s)K(s))^{-1}P(s)K(s)}_{\triangleq T_{o}(s)}]$ (34)
$E(s)$ $=Y(s)(I-T_{o}(s))=Y(s)S_{o}(s)$ (35)
If $Y(s)$ has any unstable pole $p$, the construction of $K(s)$ via (III-C) satisfies Theorem 2 and, therefore, Lemma 1 which ensure $S_{o}(p)=0$ and $S_{i}K(p)=0$. Thus, because $S_{o}(s)\in\mathcal{R}\mathcal{H}_{\infty}$ and $S_{o}(p)=0$, it follows that $E(s)=Y(s)S_{o}(s)$ is stable and bounded (i.e., $E(s)\in\mathcal{R}\mathcal{H}_{\infty}$ ). Therefore $e(t)=\mathscr{L}^{-1}\{E(s)\}$ decays exponentially, implying $\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\forall t>0.$
Similarly, since $S_{i}(s)K(s)\in\mathcal{R}\mathcal{H}_{\infty}$ and $S_{i}K(p)=0$, it follows that
$u_{inv}(t)=\mathscr{L}^{-1}\{S_{i}(s)K(s)Y(s)\}$
is also bounded. This completes the proof. ∎
Corollary 3.
Control input $u_{inv}(t)$ in Theorem 3 applies to the case where $P(s)$ has more inputs than outputs (i.e., $n_{u}>n_{y}$) but full rank.
@@
$T{\color[rgb]{0,0,0}{}_{p}}(s)\triangleq\Bigr{(}P_{p}(s)K_{p}(s)\Bigl{)}/\Bigr{(}1+P_{p}(s)K_{p}(s)\Bigl{)}$ (36)
@@
Consider $P(s)$ as defined in (11)-(12) with $\text{rank}_{\mathbb{R}(s)}(P)=n_{u}<n_{y}$ and $Z_{R}\neq\emptyset$. Then, the approximate inverse is
$\Xi_{a}^{\ell}(s)=T_{p}(s)P^{\dagger}(s).$ (37)
@@
$u_{inv}(t)=\mathscr{L}^{-1}\{T{\color[rgb]{0,0,0}{}_{p}}(s)P^{\dagger}(s)Y(s)\}.$ (38)
@@
for some $\alpha,\beta>0$. while ensuring $\|u_{inv}(t)\|_{\infty}<\infty$.
@@
Note that $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\implies U_{R1}\hat{U}_{R1}{Y}(s)={Y}(s)\implies P(s)P^{\dagger}(s)Y(s)=Y(s)$ and $P^{\dagger}(s)$, as in Corollary 2 solves (11) but with an unstable $u_{inv}(t)$ since $Z_{R}\neq\emptyset$. To get a stable $u_{inv}(t)$, we can substitute $T_{p}(s)P^{\dagger}(s)Y(s)$ into (33) as
$E(s)=Y(s)-P(s)T_{p}(s)P^{\dagger}(s)Y(s)$
and since $T_{p}(s)$ is scalar, using $P(s)P^{\dagger}(s)Y(s)=Y(s)$, and $S_{p}(s)\triangleq I-T_{p}(s)$, we get $E(s)=Y(s)S_{p}(s)$.
Assume $p$ is an unstable pole of $Y(s)$. Then since $K_{p}$ in (36) is obtained by (III-C), which gives $\begin{bmatrix}I&K_{p}\\ -P_{p}&I\end{bmatrix}^{-1}\in\mathcal{R}\mathcal{H}_{\infty}$, $P(s)$ and $P_{p}(s)$ shares the same poles, we have $S_{p}(s)\in\mathcal{RH}_{\infty}$ and $S_{p}(p)=0$ and then following the proof of Theorem 3 yields $\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\quad\forall t>0.$ Finally, Lemma 1.(1) applies on $T_{p}(s)$ so that $T_{p}(p)=0$ yielding $\|u_{inv}(t)\|_{\infty}<\infty$. This completes proof. ∎
@@
Although operators such as $S_{i}(s)K(s)$, $S_{i}(s)$, and $T_{i}(s)$ are commonly used to analyze standard closed-loop responses, here we repurpose $S_{i}(s)K(s)$ (in Theorem 3) or $T_{p}(s)$ (in Theorem 4) as a direct control action - rather than using $K(s)$ solely like a conventional loop, which departs from typical closed-loop treatments. Consequently, while alternative methods can also be used to design the $K(s)$ and therefore $S_{i}(s)K(s)$ and $T_{p}(s)$ the $\mathcal{H}_{\infty}$ framework provides valuable analytical properties. In particular, it yields an “optimal” approximate inverse within our setting.
@@
Theorem 3 and Theorem 4 provide stable approximations if the solution in (III-A) is unstable. The control input $u_{inv}(t)$ in both theorems is not unique, allowing design flexibility via (29). In Theorem 3, $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ always holds as $P(s)$ is square and full rank, while relaxing this condition, $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$, in Theorem 4 introduces a non-zero $\varphi$ in (39), as shown in (9).
All results assume a nominal system $P(s)$, but real-world complexities such as modeling errors and disturbances can lead to large errors or instabilities. These implications, along with structures over $\varphi$, will be discussed in the next section.
@@ III-D Robustness
Except for a few studies [16], to the best of our knowledge, investigations regarding stable inversion are typically conducted assuming that a complete understanding of the model is available which is often not feasible. In this section, we extend the results of Section III-C by further relaxing the constraints set forth in Theorem 1. Specifically, Theorem 5 addresses how the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ can be relaxed under uncertainties, while Corollaries 4 and 5 explore the full range of possible relaxation scenarios. For the desired output, (11)-(12) can be rewritten as
@@
where $Y_{d}(s)$ denotes the desired output. Now, to align the tracking problem in (40) to the results obtained in Sections III-B and III-C the following Assumption is useful.
Assumption 2.
The desired trajectory $y_{d}(t)\triangleq\mathscr{L}^{-1}\{Y_{d}(s)\}$ and $P(s)$ satisfies:
  • (1)
    ${\color[rgb]{0,0,0}Y_{d}(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\cap\mathcal{RH}_{\infty}}$,
  • (2)
    If $z$ is a RHP zero of $P(s)$, i.e., $z\in\mathbb{C}$ s.t. $\Re(z)>0$ and $y^{*}_{z}P(z)=0$, then $Y_{d}(z)=0$,
  • (3)
    $P(s)$ is stable or stabilized.
Now we can consider uncertainties over $P(s)$ in (40). Let the perturbed plant, $P_{\Pi}(s)$, be a member of all possible plants
$P_{\Pi}(s)\in\Pi\triangleq\left\{({I}+W_{1}\Delta_{u}{W}_{2}){P}\right\}$ (41)
where $W_{1},W_{2}$ are TFMs that characterize the spatial and frequency structure of the uncertainty, $\Delta_{u}$ denotes any unknown unstructured function with $\|\Delta_{u}\|_{\infty}<1$ [27, chapter 8.1]. Moreover, the perturbed (real) system’s zeros are
@@
Based on the conditions, exact inversion (Fig. 1(c)(i)) is valid for all system classes, including non-minimum phase, square, non-square, and singular systems, though it requires an output redefinition in the feedback path, $\bar{y}_{\Delta}(t)=({z^{+}}/{z_{d}^{+}})\ast y_{\Delta}(t)$ where $z^{+}(s)$ given by (22) and the denominator is
$z^{+}_{d}(s)=\prod_{z\in Z_{R}}(s+z)^{m_{z}}\prod_{z\in Z_{C}}(s+z)^{m_{z}}(s+{z}^{*})^{m_{{z}^{*}}}.$ (50)
Approximate inversion is applicable to non-minimum phase, full-rank systems; for square systems, the loop corresponds to Fig. 1(c)(ii), and for non-square systems, it corresponds to Fig. 1(c)(iii). Then, over the nominal system $P(s)$, the feedforward control signal $u_{ff}(t)$ is obtained by solving (III-A) (Theorem 1), (31) (Theorem 3), or (37) (Theorem 4), subject to (40). Similarly, $u_{fb}(t)$ and $u_{ff}(t)$ are obtained by following the same procedures. The next question, whether the proposed loop in Fig.1 can compensate the error caused by uncertainty, is revealed by the upcoming theorem holding P1-P2 and Assumption 2.
@@
[width=]blckdiag.pdf
Figure 1: Block diagram for closing the loop for stable inversion.
@@
Let the control inputs in Fig. 1 be defined as $u_{ff}(t)=\Xi_{i}^{\ell}(y_{d}(t))$ and $u_{fb}(t)=\Xi_{i}^{\ell}(\bar{y}_{\Delta}(t))$ and designed over $P$ in (11) with $Z_{R}\neq\emptyset$. Suppose the actual system $P_{\Pi}$ be defined as in (41) with $P_{\Pi}\neq P$. For the combined control law $u_{c}(t)=u_{ff}(t)-u_{fb}(t)$, the resulting tracking error satisfies
@@
For brevity, let us denote $W_{1}\Delta_{u}{W}_{2}$ simply by $\Delta_{u}$. Although $\operatorname*{rank}_{\mathbb{R}(s)}({P}_{\Pi})\leq\operatorname*{rank}_{\mathbb{R}(s)}({P})$, it may still occur that $\text{Im}_{\mathbb{R}(s)}(P_{\Pi})\not\subseteq\text{Im}_{\mathbb{R}(s)}(P)$ implying $\text{res}[Y_{\Pi}]=\text{res}[Y_{\Delta}]\neq 0$. Following [37, pp. 283–294], we adjust the Laplace transform integral to circumvent algebraic loops. Specifically, we define:

### v2

@@ III-C Almost Necessary and Sufficient Conditions
In this section, we propose approximate remedies for non-minimum phase systems. Specifically, by using $\mathcal{H}_{\infty}$-theory, we relax the condition $\bigl{(}Y(z)=0,\ \forall z\in Z_{R}\bigr{)}$ yielding
${(e(t)\to 0,\ \|u(t)\|_{\infty}<\infty)\iff(Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P))}$
for any unknown initial state, $\bar{x}_{0}$, of the inverse system.
Assumption 1 ensures that we can assume $P(s)$ is stable, either inherently or via stabilization, thus avoiding RHP pole/zero cancellations between the physical components on the feedforward path [24, Section 4.7.1], as is typical in stable right inversion [11, 14]. Note also that a stable (or stabilized) forward system $\Xi$, thus $P(s)$, in (6) produces an unbounded $y(t)$ if and only if $u(t)$ is unbounded. Therefore, we assume $Y(s)\in\mathcal{RH}_{\infty}$ to ensure well-behaved operation.
Now, we define a virtual loop in the sense of classical feedback structure shown by Fig.1.(a). Given the system $P(s)$ as in (11), suppose we have a virtual controller, say $K(s)\in\mathbb{R}(s)^{n_{u}\times n_{y}}$. The key transfer functions within the virtual feedback loop are defined as follows: $T_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}K(s)P(s)$; $S_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}$; and $L_{i}(s)\triangleq K(s)P(s)$. By breaking the loop at the output, we have $T_{o},S_{o},L_{o}$.
@@
$\begin{bmatrix}z(t)\\ e_{\upsilon}(t)\end{bmatrix}={\begin{bmatrix}W_{P}&W_{P}P\\ -I&-P\end{bmatrix}}\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}=G\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}$ (25)
where each scalar in $W_{P}$ is as defined in [29, Equation 22], $e_{\upsilon}(t)\triangleq r_{\upsilon}(t)-y_{\upsilon}(t)$, and $z(t)\triangleq W_{P}e(t)$. The LFT is then given by $\mathcal{F}_{l}(G,K)\triangleq{G}_{11}+{G}_{12}{K}({I}-{G}_{22}{K})^{-1}{G}_{21}$ where the $\mathcal{H}_{\infty}$ control problem involves finding $K$ such that
@@
By letting $\epsilon_{i}\to 0$, we replace all approximate integrators in $K$ with pure integrators. Moreover If $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))<\mathbb{E}[N^{2}]$, then $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))=0$ where $\mathbb{E}[N^{2}]$ denotes the expected value mean square of random noise in dB .
Remark 6.
The approximate inverse described by the following theorems can be constructed with no knowledge of $\bar{x}_{0}$.
@@
Consider $P(s)$ as defined in (11)-(12) with $\text{rank}_{\mathbb{R}(s)}(P)=r=n_{y}\leq n_{u}$, and $Z_{R}\neq\emptyset$. Define the approximate inverse $\Xi_{a}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$ with unknown $\bar{x}_{0}$ as
$\Xi_{a}^{\Gamma}(s)\;\triangleq\;S_{i}(s)K(s)\;=\;\small\left(\begin{array}[]{c|c}A_{i}^{a}&B_{i}^{a}\\ \hline\cr C_{i}^{a}&D_{i}^{a}\end{array}\right),$ (27)
@@
${\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\ \forall t>0,\text{ for some }\alpha,\beta>0.}$
@@
We decompose the control input as
$u_{inv}(t)$ $=u_{inv}^{(f)}(t)+u_{inv}^{(0)}(t)=\Xi_{a}^{\Gamma}(y(t),0)+\Xi_{a}^{\Gamma}(0,\bar{x}_{0})$
$=\mathscr{L}^{-1}\{S_{i}(s)K(s)Y(s)\}+\Xi_{a}^{\Gamma}(0,\bar{x}_{0}),$ (29)
From (10), taking the Laplace transform of $e(t)$ yields
@@
Substituting $U_{inv}(s)=S_{i}(s)K(s)Y(s)+U_{inv}^{(0)}(s)$ and noting $S_{i}(s)=\bigl{(}I+K(s)\,P(s)\bigr{)}^{-1}$, we obtain
$E(s)=Y(s)$ $-P(s)\,\bigl{(}I+K(s)\,P(s)\bigr{)}^{-1}K(s)\,Y(s)$
$-P(s)\,U_{inv}^{(0)}(s).$
Define $T_{o}(s)\triangleq\bigl{(}I+PK\bigr{)}^{-1}PK$ and $S_{o}(s)\triangleq I-T_{o}(s),$ so
$E(s)\;=\;Y(s)\,S_{o}(s)\;-\;P(s)\,U_{inv}^{(0)}(s).$
By Theorem 2, we have $S_{i}(s)K(s)\in\mathcal{RH}_{\infty}$ and $S_{o}(s)\in\mathcal{RH}_{\infty}$, implying the realization of $\Xi_{a}^{\Gamma}$ is stable. Hence
$\lim_{t\to\infty}C_{i}^{a}\,e^{A_{i}^{a}\,t}\,\bar{x}_{0}=0\implies\lim_{s\to 0}sP(s)\,U_{inv}^{(0)}(s)=0.$ (31)
Consequently, since $Y(s)\in\mathcal{RH}_{\infty}$, $\mathscr{L}^{-1}\{Y(s)\,S_{o}(s)\}$ decays exponentially, and (31), we have
$\|e(t)\|\leq\alpha\,\|e(0)\|\,e^{-\beta t},\ \forall t>0,$
for $\alpha,\beta>0$. Finally, $\mathscr{L}^{-1}\{S_{i}(s)K(s)Y(s)\}$ is bounded (since $S_{i}(s)K(s),Y(s)\in\mathcal{RH}_{\infty}$), and since $\Xi_{a}^{\Gamma}(0,\bar{x}_{0})$ is finite, so $\|u_{inv}(t)\|_{\infty}<\infty$. This completes the proof. ∎
@@
$T_{p}(s)\triangleq\Bigr{(}P_{p}(s)K_{p}(s)\Bigl{)}/\Bigr{(}1+P_{p}(s)K_{p}(s)\Bigl{)}$ (32)
@@
Consider $\text{rank}_{\mathbb{R}(s)}(P)=n_{u}<n_{y}$ and $Z_{R}\neq\emptyset$. Then, for unknown $\bar{x}_{0}$, the approximate inverse is
$\Xi_{a}^{\Gamma}(s)=T_{p}(s)P^{\dagger}(s)=\;\small\left(\begin{array}[]{c|c}A_{i}^{a}&B_{i}^{a}\\ \hline\cr C_{i}^{a}&D_{i}^{a}\end{array}\right),$ (33)
@@
$u_{inv}(t)=\mathscr{L}^{-1}\{T_{p}(s)P^{\dagger}(s)Y(s)\}+\Xi_{a}^{\Gamma}(0,\bar{x}_{0}).$ (34)
@@
for some $\alpha,\beta>0$, while ensuring $\|u_{inv}(t)\|_{\infty}<\infty$.
@@
Consider the following decomposition as
$u_{inv}(t)=u_{inv}^{(f)}(t)+u_{inv}^{(0)}(t),$ (36)
and note that $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\implies U_{R1}\hat{U}_{R1}{Y}(s)={Y}(s)\implies P(s)P^{\dagger}(s)Y(s)=Y(s)$ and $P^{\dagger}(s)$, as in Corollary 2 solves (11) but with an unstable $u_{inv}(t)$ since $Z_{R}\neq\emptyset$. To get a stable $u_{inv}(t)$, we can substitute $U_{inv}(s)=T_{p}(s)P^{\dagger}(s)Y(s)+U_{inv}^{(0)}(s)$ into (30) as
$E(s)=Y(s)-P(s)T_{p}(s)P^{\dagger}(s)Y(s)-P(s)U_{inv}^{(0)}(s)$
and since $T_{p}(s)$ is scalar, using $P(s)P^{\dagger}(s)Y(s)=Y(s)$, and $S_{p}(s)\triangleq I-T_{p}(s)$, we get
$E(s)=Y(s)S_{p}(s)-P(s)U_{inv}^{(0)}(s).$ (37)
Since $K_{p}$ in (32) is obtained by (III-C), which gives $\begin{bmatrix}I&K_{p}\\ -P_{p}&I\end{bmatrix}^{-1}\in\mathcal{R}\mathcal{H}_{\infty}$, $P(s)$ and $P_{p}(s)$ shares the same poles, we have $S_{p}(s)\in\mathcal{RH}_{\infty}$. In addition, (24) yields that If $P(s)$ has a RHP zero at $z$, then $P^{\dagger}(s)$ has a RHP-pole at $z$ and $PK(I+PK)^{-1}$, has a RHP-zero at $z$ [24], which applies on $T_{p}(s)$. Thus, $T_{p}(z)=0$ yielding $T_{p}(s)P^{\dagger}(s)Y(s)\in\mathcal{RH}_{\infty}$ Then following the proof of Theorem 3 yields $\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\quad\forall t>0.$ and $\|u_{inv}(t)\|_{\infty}<\infty$. This completes proof. ∎
@@
Although operators such as $S_{i}(s)K(s)$, $S_{i}(s)$, and $T_{i}(s)$ are commonly used to analyze standard closed-loop responses, here we repurpose $S_{i}(s)K(s)$ (in Theorem 3) or $T_{p}(s)P^{\dagger}(s)$ (in Theorem 4) as a direct control action - rather than using $K(s)$ solely like a conventional loop, which departs from typical closed-loop treatments. Consequently, while alternative methods can also be used to design the $K(s)$ and therefore $S_{i}(s)K(s)$ and $T_{p}(s)$ the $\mathcal{H}_{\infty}$ framework provides valuable analytical properties. In particular, it yields an “optimal” approximate inverse within our setting.
The next corollary provides another approximate inverse, derived from Theorem 1, that also works even with singular systems but allows for less design freedom. To achieve this, one may consider a scalar $z^{+}/z_{d}^{+}$ for the output redefinition, where
$z^{+}_{d}(s)=\prod_{z\in Z_{R}\bigcap\mathbb{R}}(s+z)^{m_{z}}\prod_{z\in Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}(s+z)^{m_{z}}(s+{z}^{*})^{m_{{z}^{*}}}$
Corollary 3.
Consider $\text{rank}_{\mathbb{R}(s)}(P)\leq\min(n_{u},n_{y})$ and $Z_{R}\neq\emptyset$ with $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$. Then, the approximate inverse is
$\Xi_{a}^{\Gamma}(s)=\Xi_{i}^{\Gamma}(s)({z^{+}(s)}/{z_{d}^{+}(s)}).$ (38)
satisfying Problem 1.b with $\varphi=0$
@@
Theorems 3, 4 and Corollary 3 each provide a stable approximation when $Y(z)\neq 0$ for any $z\in Z_{R}$. Note also that the control inputs $u_{inv}(t)$ in (28)–(34) are not unique. In Theorem 3, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ always holds because $P(s)$ is square/overactuated and full rank. However, Theorem 4 and Corollary 3 permit $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$, leading to $\varphi\neq 0$ in (9).
All results assume a nominal system $P(s)$, but modeling errors, numerical problems, and unknown forward system initial states can lead instabilities. These implications, along with structures over $\varphi$, will be discussed in the next section.
@@ III-D Robustness
Stable inversion is typically studied under the assumption of complete model knowledge including initial states $x_{0}$ which is often not feasible. In this section, we further relax the constraints in Theorem 1. Specifically, Theorem 5 addresses the condition $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$ under uncertainties, while Corollaries 4-5 explore the full range of possible relaxation scenarios. Now, the desired output can be rewritten as
@@
where bounded $Y_{d}(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\cap\mathcal{RH}_{\infty}$ denotes the desired output. For the uncertainties over $P(s)$ in (39), let the perturbed plant, $P_{\Pi}(s)$, be a member of all possible plants
$P_{\Pi}(s)=\small\left(\begin{array}[]{c|c}A_{\Pi}&B_{\Pi}\\ \hline\cr C_{\Pi}&D_{\Pi}\end{array}\right)\in\Pi\triangleq\left\{({I}+W_{1}\Delta_{u}{W}_{2}){P}\right\}$ (40)
where $W_{1},W_{2}$ are TFMs that characterize the spatial and frequency structure of the uncertainty, $\Delta_{u}$ denotes any unknown unstructured function with $\|\Delta_{u}\|_{\infty}<1$ [25, chapter 8.1]. Moreover, the perturbed (real) system’s zeros are
@@
Based on the conditions, Corollary 3 (Fig. 1(c)(i)) is valid for all system classes, including non-minimum phase, square, non-square, and singular systems, though it requires an output redefinition in the feedback path, $\bar{y}_{\Delta}(t)~{}=~{}({z^{+}}/{z_{d}^{+}})~{}\ast~{}y_{\Delta}(t)$.
The other approximate inversions are applicable to full-rank systems; for square/overactuated systems, the loop corresponds to Fig. 1(c)(ii), and for non-square (underactuated) systems, it corresponds to Fig. 1(c)(iii). Then, over the nominal system $P(s)$, the feedforward control signal $u_{ff}(t)$ is obtained by solving (III-A), (27), or (33), subject to (39). Similarly, $u_{fb}(t)$ and $u_{ff}(t)$ are obtained by following the same procedures. The next question, whether the proposed loop in Fig.1 can compensate the error caused by uncertainty, is revealed by the upcoming theorem.
@@
[width=.9]blckdiag.pdf
Figure 1: Block diagram for closing the loop in stable inversion.
@@
Consider the scheme in Fig. 1.b. Let $u_{ff}(t)=\Xi_{a}^{\Gamma}(y_{d}(t),\bar{x}_{0_{1}})$ and $u_{fb}(t)=\Xi_{a}^{\Gamma}(\bar{y}_{\Delta}(t),\bar{x}_{0_{2}})$ be designed over the $P(s)$ with $Z_{R}\neq\emptyset$. Suppose the actual system $P_{\Pi}$ (with $P_{\Pi}\neq P$ and unknown $x_{0}$) is as in (40). Then, under $u_{c}(t)=u_{ff}(t)-u_{fb}(t)$ the tracking error satisfies
@@
For brevity, let us denote $W_{1}\Delta_{u}{W}_{2}$ simply by $\Delta_{u}$, and note that $u_{c}\triangleq u_{c}^{(f)}+u_{c}^{(0)}=u^{(f)}_{ff}+u^{(0)}_{ff}-u^{(f)}_{fb}-u^{(0)}_{fb}$. Although $\operatorname*{rank}_{\mathbb{R}(s)}({P}_{\Pi})\leq\operatorname*{rank}_{\mathbb{R}(s)}({P})$, it may still occur that $\text{Im}_{\mathbb{R}(s)}(P_{\Pi})\not\subseteq\text{Im}_{\mathbb{R}(s)}(P)$ implying $\text{res}[Y_{\Pi}]=\text{res}[Y_{\Delta}]\neq 0$. Without losing generality, we can re-define the bounds of integration as [32, p.283-294]